OpenStudy (anonymous):

Need answers to my question giving medal and fan!! Write and equation in point slope form for the line with the given slope and through the given point. 1. (0,4);m=3 2. (-2,0);m=-4 3. (5,-3);m=-3/5 4. (-1,-9);m=-2/3 5. (-6,4);m=-3/5 6. (7,11);m=1/2

OpenStudy (anonymous):

Point slope form is: y – y1 = m(x – x1) So for third example: y + 3 = (-3/5)(x - 5) Can you do the rest?

OpenStudy (anonymous):

ill try

OpenStudy (anonymous):

ok

OpenStudy (mathmale):

finished solving this problem? How about an update?